Riley Company provides the following information for Material A:

Actual quantity purchased and used 6,500 lbs.

Standard quantity allowed 6,000 lbs.

Actual price $3.80

Standard price $4.00

A. What was the direct material quantity variance for Material A?

B. What was the direct material price variance for Material A?
Answer #1

Materials Quantity Variance = (SQ-AQ) * SP

= (6000 - 6500) * 4

= 2000 Unfavorable

Materials Price Variance = (SP-AP) * AQ

= (4 - 3.80) * 6500

= 1300 Favourable